A tennis player receives a shot with the ball (0.0600 kg) traveling horizontally at 58.0 m/s and returns the shot with the ball traveling horizontally at 36.8 m/s in the opposite direction. (Take the direction of the ball's final velocity (toward the net) to be the +x-direction.)


(a) What is the impulse delivered to the ball by the racket?

magnitude ___ N · s


(b) What work does the racket do on the ball? (Indicate the direction with the sign of your answer.)

 ___J


1
Expert's answer
2021-10-02T10:46:55-0400

a) Impulse= Change of momentum

"J = 0.06(36.8 -(-58.0)) = 5.688 \; N \cdot s"

b) Work done by the racket,

"W = \frac{1}{2}m(v^2_2-v^2_1) \\

= 0.5 \times 0.06(36.8^2 -58.0^2) \\

= 0.03(1354.24 -3364) \\

= -60.293 \;J"